OpenStudy (anonymous):

The area of a triangular section of a store cannot be more than 910 sq. ft. If the height of the triangular section is 52 feet, what should the base be?

OpenStudy (anonymous):

Area = 1/2 x base x height Let the base be b 910 = 1/2 x b x 52 910 = 26b 35 = b Therefore the base is 35 ft

OpenStudy (anonymous):

hope this helps! :)

OpenStudy (anonymous):

Got it thank you